OpenStudy (anonymous):

simplify 3(n+1)! / 5n! without a calculator

OpenStudy (vishweshshrimali5):

(n+1)! = (n+1) n! so \[\cfrac{3(n+1)!}{5n!} = \cfrac{3(n+1)n!}{5n!}\] simplify this

OpenStudy (anonymous):

wait why does (n+1)! = (n+1)n! ?

OpenStudy (anonymous):